OpenStudy (anonymous):

simpify 7a - (3a + 4)

OpenStudy (anonymous):

7a - (3a + 4) 7a - 3a - 4 4a -4 = 0 +4 +4 4a = 4 /4 /4 a = 1

OpenStudy (anonymous):

thank you